the length of a rectangle is four times its width. if the area of the rectangle is 400 m2 find the perimeter
Answer by satyareddy22(84)

Let l be length,b be the breadth.
l=4b (given)
Area = 400 m2
But area of rectangle = l * b
l * b = 400
(4b)b=400
4b^2=400
b^2=400/4=100
b=sq(100)=10
b=10
l=4b=4(10)=40
l=40
Perimeter = 2(l+b)=2(10+40)=2(50)=100